X Company has two production departments, A and B. The following is budgeted information for all of its products in 2019, and actual information for one of them, Product X:

Department A All Products Product X
   Overhead $2,000,000 --   
   Direct labor hours 40,000 894  
   Machine hours 100,000 1,000  
   Units produced 54,000 700  
Department B
   Overhead $2,100,000 --   
   Direct labor hours 60,000 202  
   Machine hours 140,000 890  
   Units produced 46,000 700  

Using a single-driver allocation system, with machine hours as the driver, how much overhead was allocated to Product X [round overhead rate(s) to two decimal places]
Answer #1

Solution:

Predetermined overhead rate = Estimated overhead / Estimated machine hours

= ($2,000,000 + $2,100,000) / (100000+140000) = $17.08 per machine hour

Overhead allocated to Product X = (1000 + 890) * $17.08 = $32,281.20